Suppose that the quantity of coffee demanded declines from 206,000 gallons per week to 194,000 gallons per week as a consequence of a 10 percent increase in the price of coffee drinks. The price elasticity of demand is a. 1.66. b. 6.00. c. 0.60. d. 1.40. Show the calculation.


1
Expert's answer
2022-03-08T12:54:47-0500

"PED=\frac{\% change in quantity demanded}{\% change in price}"


% change in quantity

"=\frac{-12000}{206000}×100=-5.825"


%change in price is 10

Therefore PED is

"=\frac{5.825}{10}=0.5825"